OpenStudy (dragontears1128):

LAW ENFORCEMENT: A police accident investigator can use the formula S=2√5L to estimate the speed s of a car in miles per hour based on the length l in feet of the skid marks it left. How fast was a car traveling that left skid marks 129 feet long? Round your answer to the nearest miles per hour

OpenStudy (kl0723):

ok, so here you have another function to evaluate, it's just called s, and s = \[2\sqrt{5L}\]

OpenStudy (kl0723):

so this evaluates the speed, given the length of the skid marks when the car started to brake, so basically evaluate s when L=129 feet in this case

OpenStudy (kl0723):

\[s = 2\sqrt{5(129)}\] plug in the value of L

OpenStudy (kl0723):

what do you get?

OpenStudy (dragontears1128):

do i simplify it

OpenStudy (kl0723):

yes, you can use a calculator

OpenStudy (kl0723):

5 times 129 = 645

OpenStudy (dragontears1128):

thats what i got

OpenStudy (dragontears1128):

when i do it

OpenStudy (dragontears1128):

done

OpenStudy (kl0723):

now \[\sqrt{645}\] yields 25.3968

OpenStudy (dragontears1128):

yep

OpenStudy (kl0723):

and 2(25.3968) = 50.7937

OpenStudy (kl0723):

but the problem says round to the nearest mile per hour, so you can say that the car was traveling at 51 miles per hour

OpenStudy (kl0723):

did my answer just get deleted? lol

OpenStudy (dragontears1128):

im doing a review for my exam and nope lol

OpenStudy (kl0723):

oops it just didn't refresh haha i see it now

OpenStudy (kl0723):

yeah, that was like the previous problems you've done but this time around is just a word problem, make sure you read it and identify the given values so you can compute the function easily

OpenStudy (dragontears1128):

im just not good with word problems

OpenStudy (dragontears1128):

tbh

OpenStudy (kl0723):

they can be tricky at times

OpenStudy (dragontears1128):

thank you again